# A call for frugal modelling: two case studies involving molecular spin dynamics

### v2

@@ Abstract
An additional question, beyond the model we choose to solve, is how we actually solve it. The ease of use of programming languages has been being increasing over time, with a rising popularity of high-level computing environments. These provide interactive exploratory environments that make language features and libraries immediately available to scientists who can use them to explore a problem domain. This contrasts with the classical edit/compile/run cycle of C or Fortran programming, which typically requires separate computation and post-processing/visualization steps. 48 Additionally, the popularity of high-level languages means solutions to syntax errors and general advice can be found on the Internet. However, high-level programming often comes at a cost in resource use.
@@
Let us take an established physical model and deal only with the implementation choice. We now explore an already published program that models molecular nano-magnets and their application as probabilistic bits, called STOSS, 21 and test a different implementation of the same algorithm.
@@
Read input data from EXCEL file |
user configurations | 0.2346 | 0.2080 |
Read system characteristics from EXCEL file |
read_data | 0.0725 | 0.0094 |
Calculate magnetic relaxation |
mag_relaxation | 0.0021 | 0.0010 |
Calculate probabilities of each spin to flip in the 1st p-bit |
Bolztmann_distribution | 0.0010 | 0.0010 |
Iteration process (”for” loop) for the 1st p-bit |
changeable_field | 305 | 16800 |
Calculate probabilities of each spin to flip in the 2nd p-bit |
Bolztmann_distribution | 0.0473 | 0.2630 |
Iteration process (”for” loop) for the 2nd p-bit |
changeable_field | 275 | 16900 |
Average p-bits states over time |
mean_matrix_state | 0.0460 | 0.0036 |
Association analysis between both p-bits |
association | 0.1484 | 2.3946 |
Plotting results |
plotting | 0.7876 | 0.4762 |
Total time |
581 | 33700 | |
STOSS consists in a custom implementation of a Markov Chain Monte Carlo algorithm for each of the independent particles (in this case, effective spins ). The relative Markov chain probabilities for the spin flip between ground and excited spin states correspond to the relative Boltzmann populations of the two effective spin states , . Each computational step has an associated natural time duration that is derived from parameterized average spin dynamics, thus the model allows one to follow independent time trajectories. There are three main scenarios studied using experimental data for comparison, and all the details could be found in the Supporting Information File of Gutiérrez-Finol et al 21.
Each implementation of the same program, even if following a given algorithm as closely as possible, have distinct costs, in terms of memory use, runtime and energy consumption. This has often been analyzed in particular for implementations employing different programnming languages. 49 In the case of Matlab vs Python a major difference arises in ”for” loops, where Matlab is faster than Python. More generally, M language has a strongly typed syntax, often resulting in a improvements in memory usage and processing time. Identifying the type of each variable at compile-time allows the compiler to optimize the code, saving time and being able to use the minimum amount of memory. That being said, we are not claiming here that the behaviour we present here is univocally associated with implementing the model in python vs in M language, since many different approaches are always possible even within a given language and algorithm.
@@
We performed the calculations on the third case of study in a lanthanide-based, molecular spin p-bit network 21 which corresponds to the longest computation in the paper. Having in mind the original case, we implemented the simulation maintaining the same conditions in both scenarios, therefore, we explored a 2-p-bit architecture where each p-bit is constituted by the collective signal of magnetic molecules which evolve freely, with the molecules corresponding to the first p-bit evolve in absence of a magnetic field and the one corresponding to the second p-bit evolve in presence of a magnetic field determined by the state of the first p-bit. The program is divided in seven functions where each one accomplishes a specific task, two functions are reused for the calculation of the probability of each p-bit to change its state. Table 2. presents the time speed processing for each function and as it could suggest by far the largest difference is from the ”for” loops when the program iterates over each spin at each time step.
@@
In this case the overall total runtime is rather short in any case, as are the associated emissions. However, should one employ STOSS to fit experimental data, as was done recently, 36 one would be tempted to explore a wide parameter range, and that would give rise to a larger carbon footprint if the more expensive implementation was used. A good approach here would be, as soon as the expected calculation time veers into the hundreds or thousands of hours, start thinking about a less expensive implementation. Of course it can be even better to optimize the model rather than just the implementation, and we are indeed working on that.
Conclusion
@@
We presented here two particular examples illustrating an extremely common situation: when confronted to a calculation with a large carbon footprint, one can often choose either to solve a different model or to solve the same model via a different implementation, and obtain significant savings in carbon footprint without making any significant sacrifices in scientific yield. Currently this is mostly overlooked, and if anything the most expensive methods tend to enjoy a higher prestige and are considered more trustworthy. Often, this means that wasteful methods allow for easier or better publishing venues. In our case, the research presented in Liu et al 35 was originally submitted with the frugal method, but during the refereeing process we were requested to switch to the method that is in principle more exact but which we show here is actually wasteful in this case. As a community we need to do better, and the main factor is choosing affordable models to solve problems, with a minor but sizeable contribution of finding an inexpensive way to implement these models. Crucially, we need to beware of the Jevons’ paradox or rebound effect, 37 meaning a more efficient method, if not coupled to resource consciousness, by itself leads to an increased usage which often overshoots the savings. Thus, we call herein for ”frugality” rather than for ”efficiency”. More generally, just as we consider the ethical repercussions of animal experimentation, or dealing with patient data, eventually we will need to include the risk of carbon footprint wastefulness as an ethical concern in research.
@@
For the second section of this paper, the original code (python version) named STOSS (for STOchastic Spin Simulator) is available at https://github.com/gerlizg/STOSS. The new version of the program using Matlab and the instructions to reproduce all the results could be found in the Supporting Information Section S2 and at https://github.com/gerlizg/STOSS-MATLAB.

### v3

@@ Abstract
An additional question, beyond the model we choose to solve, is how we actually solve it. The ease of use of programming languages has been being increasing over time, with a rising popularity of high-level computing environments. These provide interactive exploratory environments that make language features and libraries immediately available to scientists who can use them to explore a problem domain. This contrasts with the classical edit/compile/run cycle of C or Fortran programming, which typically requires separate computation and post-processing/visualization steps. 59 Additionally, the popularity of high-level languages means solutions to syntax errors and general advice can be found on the Internet. High-level programming often comes at a cost in resource, but at the same time, these languages are designed to be easy for humans to read and write, allowing developers to focus more on problem-solving rather than worrying about low-level details like memory management or hardware specifics. Nevertheless, they can support green coding practices by facilitating efficient coding techniques and resource management. While high-level languages are often not as efficient as low-level ones (e.g., C or assembly), they encourage code practices that can still lead to efficient performance through optimized libraries, efficient algorithms and data structures, contributing to lower energy consumption.
@@
Let us take an established physical model and deal only with the implementation choice. We now explore an already published program that models molecular nano-magnets and their application as probabilistic bits, called STOSS, 33 and test a different implementation of the same algorithm.
@@
Read input data from EXCEL file | user configurations | 0.2346 | 0.2080 |
Read system characteristics from EXCEL file | read_data | 0.0725 | 0.0094 |
Calculate magnetic relaxation | mag_relaxation | 0.0021 | 0.0010 |
Calculate probabilities of each spin to flip in the 1st p-bit | Bolztmann_distribution | 0.0010 | 0.0010 |
Iteration process (”for” loop) for the 1st p-bit | changeable_field | 305 | 16800 |
Calculate probabilities of each spin to flip in the 2nd p-bit | Bolztmann_distribution | 0.0473 | 0.2630 |
Iteration process (”for” loop) for the 2nd p-bit | changeable_field | 275 | 16900 |
Average p-bits states over time | mean_matrix_state | 0.0460 | 0.0036 |
Association analysis between both p-bits | association | 0.1484 | 2.3946 |
Plotting results | plotting | 0.7876 | 0.4762 |
Total time | 581 | 33700 |
STOSS consists in a custom implementation of a Markov Chain Monte Carlo algorithm for each of the independent particles (in this case, effective spins ). The relative Markov chain probabilities for the spin flip between ground and excited spin states correspond to the relative Boltzmann populations of the two effective spin states , . Each computational step has an associated natural time duration that is derived from parameterized average spin dynamics, thus the model allows one to follow independent time trajectories. There are three main scenarios studied using experimental data for comparison, and all the details could be found in the Supporting Information File of Gutiérrez-Finol et al 33.
Each implementation of the same program, even if following a given algorithm as closely as possible, have distinct costs, in terms of memory use, runtime and energy consumption. This has often been analyzed in particular for implementations employing different programnming languages. 60 In the case of Matlab vs Python a major difference arises in ”for” loops, where Matlab is faster than Python. More generally, M language has a strongly typed syntax, often resulting in a improvements in memory usage and processing time. Identifying the type of each variable at compile-time allows the compiler to optimize the code, saving time and being able to use the minimum amount of memory. That being said, we are not claiming here that the behaviour we present here is univocally associated with implementing the model in python vs in M language, since many different approaches are always possible even within a given language and algorithm.
@@
We performed the calculations on the third case of study in a lanthanide-based, molecular spin p-bit network 33 which corresponds to the longest computation in the paper. Having in mind the original case, we implemented the simulation maintaining the same conditions in both scenarios, therefore, we explored a 2-p-bit architecture where each p-bit is constituted by the collective signal of magnetic molecules which evolve freely, with the molecules corresponding to the first p-bit evolve in absence of a magnetic field and the one corresponding to the second p-bit evolve in presence of a magnetic field determined by the state of the first p-bit. The program is divided in seven functions where each one accomplishes a specific task, two functions are reused for the calculation of the probability of each p-bit to change its state. Table 3 presents the time speed processing for each function and as it could suggest by far the largest difference is from the ”for” loops when the program iterates over each spin at each time step.
@@
In this case the overall total runtime is rather short in any case, as are the associated emissions. However, should one employ STOSS to fit experimental data, as was done recently, 48 one would be tempted to explore a wide parameter range, and that would give rise to a larger carbon footprint if the more expensive implementation was used. A good approach here would be, as soon as the expected calculation time veers into the hundreds or thousands of hours, start thinking about a less expensive implementation. Of course it can be even better to optimize the model rather than just the implementation, and we are indeed working on that. Additionally, we now present an updated version of the program, implementing several optimizations to enhance performance. By leveraging the vectorized operations and built-in functions available in the MATLAB programming language, we significantly reduced computational time. These optimizations allow for more efficient data handling and processing, minimizing the need for iterative loops. As a result, the program is now capable of delivering faster results without compromising accuracy, offering an overall improvement in both speed and functionality.
Conclusion and context
The most general question we are dealing with here is how to take into account the goal of environmental sustainability when doing science; in the case of chemistry this is the field of green chemistry. A closely related question is how to balance environmental sustainability and scientific advancement. Our proposal is to aim for a rational cost-benefit analysis, and a prerrequisite for this is being aware of the costs. Only after researchers, research groups and research institutions are quantitatively aware of how their carbon footprints derives from their choices, we will be able to rationally reduce emissions.
Quantifying CO2 footprint in chemistry laboratories is far from a trivial task and requires and exhaustive life cycle assessment (LCA). We should account for everything when assessing an LCA: materials acquisition and input, manufacturing and production, packaging and distribution, product use, disposal and recycling).61 Consequently, in the framework of green chemistry, we urge to implement objectives of sustainability also into computational chemistry research, not only to design safer chemicals and processes to avoid physical experiments, but also to reduce energy consumption and minimize waste. Just as we aim to prevent and minimize when designing a chemical reaction, we should also standardize preventive practices during computational experiments.
Here we need to be aware of the urgency of the climate crisis and thus avoid relying on unproven solutions and risky bets on techno-optimistic futures. In particular, in the field of theoretical and quantum chemistry, it would be irresponsible to wait until some supposedly energy efficient quantum computers start solving practical problems. Presently, both the operation of quantum computers and the simulation of quantum circuits carry a considerable carbon footprint and are not being used to avoid the footprint of conventional supercomputers.62, 63 The same can be said of AI-based solutions, which may well bring increased productivity but have so far substantially increased the overall computing carbon footprint, mainly through the training cost of the models, with no clear path to net reduction.
We presented here two particular examples from the field of computational materials science illustrating an extremely common situation: when confronted to a calculation with a large carbon footprint, one can often choose either to solve a different model or to solve the same model via a different implementation, and obtain significant savings in carbon footprint without making any significant sacrifices in scientific yield. Currently this is mostly overlooked, and if anything the most expensive methods tend to enjoy a higher prestige and are considered more trustworthy. Often, this means that wasteful methods allow for easier or better publishing venues.Indeed, this is a general problem rather than exclusive for computation: a wasteful excess of experimental techniques is rarely if ever seen as a problem from the publishing perspective, and only a matter of money.
@@
In our case, the research presented in Liu et al 47 was originally submitted with the frugal method, but during the refereeing process we were requested to switch to the method that is in principle more exact but which we show here is actually wasteful in this case. As a community we need to do better, and the main factor is choosing affordable models to solve problems, with a minor but sizeable contribution of finding an inexpensive way to implement these models. Crucially, we need to beware of the Jevons’ paradox or rebound effect, 26 meaning a more efficient method, if not coupled to resource consciousness, by itself leads to an increased usage which often overshoots the savings. This can be seen as a particular case of the necessary attitudinal change in all of chemistry, 64. Thus, we call herein for ”frugality” rather than for ”efficiency”. More generally, just as we consider the ethical repercussions of animal experimentation, or dealing with patient data, eventually we will need to include the risk of carbon footprint wastefulness as an ethical concern in research.
@@
For the second section of this paper, the original code (python version) named STOSS (for STOchastic Spin Simulator) is available at https://github.com/gerlizg/STOSS. The new version of the program using Matlab and the instructions to reproduce all the results could be found in the Supporting Information Section S2 and at https://github.com/gerlizg/STOSS-MATLAB. Finally, the optimized version of the simulator could be found at https://github.com/gerlizg/STOSS-Optimized-Matlab-Version.